OpenStudy (anonymous):

How do you find the slope of the following functions at the given points?: f(x)=-1/2+7/5x^3 at the point (3/5, 2) y=(2x+1)^2 at the point (0,1) f(x)=3(5-x)^2 at the point (5,0) f(theta)=4sin(theta)-(theta) at the point (0,0) g(t)=2+3cos(t) at the point (pi, -1)

OpenStudy (anonymous):

Find first derivative, then plug in the x (or theta or t) value from the point given.

OpenStudy (anonymous):

I'm having trouble finding the derivative for some for these. I'll work out one of the problems and tell me what I did wrong. f(x)=3(5-x)^2 for the point (5,0) f(x)=3(25-10x+x^2) f(x)=75-30x+3x^2 f'(x)=75-30(x^0)+3(2x) f'(x)=75-30+6x f'(5)=30 The answer key and the calculator says that the answer is 0 instead of 30. Where did I go wrong?

OpenStudy (anonymous):

When you take the derivative of a constant, e.g. that 75 there, it becomes zero.

OpenStudy (anonymous):

You can also differentiate via chain rule so you don't have to simplify the expression first.

OpenStudy (anonymous):

\[f(x)=3(5-x)^2 \rightarrow f'(x)=-1 \cdot 2 \cdot 3 \cdot (5-x)^1 \rightarrow -6(5-x)\] At x=5, f'(5)=0.

OpenStudy (anonymous):

I got lost at the second step. Can you tell me what you did?

OpenStudy (anonymous):

Sure, the chain rule says that when you differentiate a function of a function ((5-x)^2 is a function of 5-x, which itself is a function of x.) you multiply the derivative of the inside function by the derivative of the outside function.

OpenStudy (anonymous):

For 3(5-x)^2, first multiply by the derivative of the inside function (the derivative of 5-x is -1), then take the derivative of the squaring function which is 2.

OpenStudy (anonymous):

Altogether, f'(x) = -1*2*3*(5-x) = -6(5-x). Don't bother simplifying to 6x-30. Plugging 5 into (5-x) makes it obvious that the answer is zero.

OpenStudy (anonymous):

Is there another one you're having trouble with?

OpenStudy (anonymous):

Yes. The first problem I posted and one of the trig problems. (Thanks for the explanation on the third problem. I think I understand it now!)

OpenStudy (anonymous):

Show me what you did for the first one and I'll try the f(theta) one.

OpenStudy (anonymous):

The derivative of 4sin(Θ) - Θ = 4cos(Θ) - 1. At Θ=0, f' = 3.

OpenStudy (anonymous):

Okay, so this is what I did: f(x)=-1/2+7/5x^3 at the point (0, -1/2) [sorry, I wrote the point down wrong in my question] f(x)=-1/2+7/5(3x^2) f'(x)=-1/2+(21/5)x^2 f'(0)=0 Is this the right way to solve this one?

OpenStudy (anonymous):

Everything is fine except remember that the derivative of a constant is zero, so f'(x)=21x^2/5.

OpenStudy (anonymous):

So, wait how are denominators and numerators switched?

OpenStudy (anonymous):

Sorry. Took a food break. Denominators and numerators? What got switched?

OpenStudy (anonymous):

So I had it as f'(x)=-1/2+(21/5)x^2 Then you put it as f'(x)=21x^2/5.

OpenStudy (anonymous):

yep, 21-over-5. Same as you. I just wrote it slightly differently.

OpenStudy (anonymous):

\[\frac{21}{5}x^2 = \frac{21x^2}{5}\]

OpenStudy (anonymous):

Oh. Okay. I have another question over finding the derivative of this problem: \[f(x)=\frac{x ^{3}-3x^{2}+4}{ x^2 }\] I have no clue how to do this one.

OpenStudy (anonymous):

i. Divide out x^2 so you get f(x) = x - 3 + 4/(x^2) Rewriting the last term, f(x) = \[x - 3 + 4x ^{-2}\] Finding the derivative should be more manageable now